OpenStudy (vortish):

solve by factoring 5x^2+6x-63=0

OpenStudy (anonymous):

x = -21/5, 3

OpenStudy (anonymous):

\[5x ^{2}+6x-63=(5x+21)(x-3)\] 5x+21=0 5x=-21 x=-21/5 x-3=0 x=3